Status Report

NASA Space Station On-Orbit Status 29 January 2005

By SpaceRef Editor
January 29, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 29 January 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below. Saturday — first weekend rest day for the crew.

Finishing up on post-EVA cleanup activities, FE Salizhan Sharipov started the discharge process on the second of two Orlan 825M3 battery units, terminating it later at ~2:25pm EST.

CDR/SO Chiao and Sharipov performed the regular weekly 3-hr. task of thorough station cleaning. [“Uborka”, done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

Dr. Chiao’s selection for today’s “Saturday Science” program was CBOSS-FDI (Cellular Biotechnology Support Systems-Fluid Dynamics Investigation) Tissue Culture Module (TCM) and syringe bubble removal. The techniques used to remove bubbles from the TCM and syringe will be critical to future CBOSS on-orbit experiments. [Leroy started with a review of pertinent training material, concentrating on three techniques (TCM Resuspension, TCM Arc Swing, and Syringe Arc Swing) and also reviewed a synopsis of bubble removal procedures. Supported by a teleconference with the Principal Investigator and his NIH (National Institutes of Health) team at 11:30am, the Science Officer then performed the activities. Bubbles were an unexpected feature in TCM on ISS, discovered in the high quality images obtained by Ed Lu and Mike Foale during previous increments. Bubble removal is now a new focus since bubbles are known to kill cells. Mike Foale spoke of observing two different ranges of bubble size with different behavior: large and small (almost foamy). Experimenters need to make a decision as to which bubble removal method to use in future cell-based experiments.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Sharipov did the routine maintenance of the Service Module (SM) environment control & life support system (ECLSS/SOZh), including ASU toilet facilities and the weekly inspection of the air/liquid condensate separator apparatus (BRPK).

The crew performed their regular daily 2.5-hr. physical exercise program on TVIS, RED exerciser, CEVIS cycle and the VELO cycle with bungee cord load trainer. [Sharipov’s daily protocol prescribes a 1.5-hr structured set on the treadmill and one hour on VELO (today: Day 3 of a new set).]

The CDR then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

Working off the discretionary Russian task list, Salizhan prepared payload service data files of the new commercial “RokvISS” experiment from yesterday’s ground-controlled checkout. [These data, along with log files, were transferred from the BSPN payload matching unit (Server) to the ISS Wiener laptop for subsequent downlink via OCA comm as a ZIP file, for ground analysis of the German payload.]

Yesterday’s sampling and testing of ITCS (Internal Thermal Control System) coolant by Leroy Chiao for presence of NH3 (ammonia) came up with nominal results (0.25 ppm, parts per million), indicating no ammonia leak. [NH3 is the coolant used by the EETCS (Early External Thermal Control System). The ITCS accepts heat from ISS internal thermal loads and transfers it to the EETCS through the IFHX (Interface Heat Exchanger). The EETCS then rejects the heat to space via radiators. The EETCS operates at a higher pressure than the ITCS, and periodic tests are intended to determine whether or not there is ammonia leakage through the IFHX.]

During yesterday’s regular weekly PCS (Portable Computer System) reboots by Chiao, the HDD (hard disk drive) of the SM PCS was found to be corrupted. It was replaced with HDD #6026, which was then loaded with the recently uplinked BGAD (Beta Gimbal Assembly Display) patch.

The Elektron oxygen generator remains off, per post-EVA plan. There are currently no reported problems with the unit.

Total propellants currently available in FGB, SM and Progress 16: 3761 kg (8292 lbs).

Weekly Science Update (Expedition Ten — 13th):

GASMAP: Nothing new.

Human Research Facility/Workstation (HRF WS): Continuing.

Advanced Ultrasound (ADUM): Nothing new.

Hand Posture Analyzer (HPA: Nothing new.

In-Space Soldering Investigation (ISSI): Operations are complete.

Pore Formation and Mobility Investigation (PFMI): Nothing new.

Space Acceleration Measurement System (SAMS): Nothing new.

Microgravity Acceleration Measurement System (MAMS): Nothing new.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS): Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE): Nothing new.

Binary Colloidal Alloy Test-3 (BCAT-3): Nothing new.

Renal Stone (RS): Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES): Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT): Nothing new.

Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): Nothing new.

Biotechnology Specimen Temperature Controller (BSTC): Nothing new.

Yeast Group Activation Packs (Yeast GAP): Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM): A new camera, laptop, software combination will be used for EarthKAM ops this week, 2/1-2/5. A record number of schools are enrolled. Organizers are expecting their “best and busiest week ever”.

Earth Science Toward Exploration Research (ESTER): Nothing new.

Serial Network Flow Monitor (SNFM): Nothing new.

Fluid Merging Viscosity Measurement (FMVM): Nothing new.

Space Experiment Module (SEM): The SEM team thanked the crew for their “great work so far”. The students are enjoying the photos from the first set of operations this month, as well as the video interview with Central Park Middle School. Next activity will be a SEM video introduction on 2/2, a high point because it will be the first time in the long history of SEM that such footage has been taken.

Viscous Liquid Foam–Bulk Metallic Glass (Foam): Nothing new.

Effects of Prolonged Space Flight on Human Skeletal Muscle (BIOPSY): Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.

Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.

Miscible Fluids in Microgravity (MFMG): Nothing new.

Educational Payload Operations (EPO): Nothing new.

Capillary Flow Experiment (CFE): Nothing new.

Crew Earth Observations (CEO): The ground has received and continues to review over 8,500 CEO images from this Increment to date. A nicely detailed long lens view of the Parana River floodplain in Argentina will be posted on Earth Observatory this week. The 800mm image details the complex and poorly understood landforms and vegetation patterns that result when a large river, such as the Parana, abandons a major channel in its floodplain. “Nice shot! The quality and composition of your imagery is masterful.”

Today’s CEO (crew earth observations) photo targets, no longer limited in the current LVLH attitude, were Internal Waves, Caribbean Sea (this overpass provided an opportunity for internal wave photography. Looking to the right of track, and to the south of Hispaniola, for the sunglint point), and Johnston Island reef, Central Pacific (weather was predicted to be mostly clear over this reef. Detailed imagery of the island and fringing reefs are useful for updating geographic maps and assessing reef health).

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Key Events:

  • Progress 16 (16P) undocking & destructive reentry — 2/27/05;
  • Progress 17 (17P) launch — 2/28/05;
  • EVA-13 — 3/25/05;
  • Soyuz 10 (10S) launch — 4/15/05 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
  • Soyuz 9 (9S) undock — 4/25/05 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
  • Progress 18 (18P) launch — 6/10/05;
  • Progress 19 (19P) launch — 8/24/05;
  • Soyuz 11 (11S) launch — 9/27/05.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit (as of this morning, 10:50am EST [= epoch]):

  • Mean altitude — 357.7 km
  • Apogee height — 364.3 km
  • Perigee height — 351.0 km
  • Period — 91.69 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0009899
  • Solar Beta Angle — -58.1 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 60 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 35390

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.